# Grokking Group Multiplication with Cosets

### v1

@@ 5.2 Conjugate Subgroup Circuit
All four ways to multiply two cosets of $A_{n}$ are well-defined. For each of the four options (even-even, odd-even, etc…) we know which coset of $A_{n}$ the product will be in, but no other subgroup of $S_{n}$ has this property. What they have are other conjugate subgroups. The model learns to identify those subgroups using the fact that the coset multiplication is well-defined for specific pairs of conjugate subgroups. In table 1 we provide the coset multiplication table for two subgroups of $S_{5}$ that we will refer to as $H_{1}$ and $H_{5}$. Both are isomorphic to $S_{4}$, $H_{1}$ fixes $1$ and $H_{5}$ fixes $5$. They are related by conjugation. The transposition $\tau_{(15)}$ sends one subgroup to the other: $\tau_{(15)}H_{1}\tau_{(15)}=H_{5}$ and $\tau_{(15)}H_{5}\tau_{(15)}=H_{1}$.
Far from being fully defined for every combination, every right coset of $H_{5}$ has precisely one paired left coset of $H_{1}$ that produces the coset $H_{1}\tau_{(15)}$. $H_{1}\tau_{(15)}$ has 24 elements, but the because of the double coset structure of
Despite this being more complicated for us to reason about than the $A_{5}$ case, the neural network treats these cases the same. There are pairs of subgroups that have a predictable multiplication, for a specific neuron the model learns features so that the pre-activations of the predictable products precisely cancel out to zero, and when that neuron fires the models knows that the product is not in the coset that neuron is specialized to predict.
$H_{1}$ $\tau_{(12)}H_{1}$ $\tau_{(13)}H_{1}$ $\tau_{(14)}H_{1}$ $\tau_{(15)}H_{1}$
$H_{5}$ $\tau_{(15)}H_{1}$
$H_{5}\tau_{(15)}$ $\tau_{(15)}H_{1}$
$H_{5}\tau_{(25)}$ $\tau_{(15)}H_{1}$
$H_{5}\tau_{(35)}$ $\tau_{(15)}H_{1}$
$H_{5}\tau_{(45)}$ $\tau_{(15)}H_{1}$
Table 1: Coset multiplication table for $H_{1},H_{5}$ where $H_{i}\coloneqq\{\sigma\in S_{5}|\;\sigma(i)=i\}$. The coset $\tau_{(15)}H_{1}$ corresponds to all of the permutations in $S_{5}$ with $1$ in the fifth position. If the space is left blank, then multiplication of those two cosets creates every element of $S_{5}$ not in $\tau_{(15)}H_{1}$.
@@ 5.3 Decoding Permutations with Coset Membership
Figure 4: An example of an $S_{4}<S_{5}$ coset circuit bit pattern for the permutation (5 4 3 2 1)54321(5\;4\;3\;2\;1)( 5 4 3 2 1 ). The value of each position of the permutation is encoded in five coset circuits, twenty-five in total. Highlighted is the code for the third position. The value is $3$ because three is the only circuit for that position that did not fire. Each individual circuit for a position-value pair (i.e. a single square in the diagram) will consist of as few as one or as many as fifteen neurons dedicated. The median number of neurons involved in a single coset circuit depends on the subgroup, but is around four.
Figure 5: The effect of combining coset circuits. These show the contributions of one, two, and all four $S_{4}$ coset circuits to the logits of each permutation when the target permutation is not in $S_{4}$ (when the target is in $S_{4}$ then none of these neurons fire and their contributions to the logits is identically zero).
The unembedding layer receives as input a vector of activations from that encode the coset membership of of the product $\sigma_{l}\sigma_{r}$. Each element of this vector can best be thought of as an on/off switch keyed to a specific coset. If the activation is positive then the product $\sigma_{l}\sigma_{r}$ is certainly not in that coset, but if it is zero it is only probably in the coset. Putting aside this ambiguity for a moment, we can see that with enough of these circuits the model can uniquely encode every permutation $\sigma\in S_{n}$ and perfectly solve the group multiplication.
@@
The actual single-neuron coset circuits are not perfect on/off switches. Because of the ReLU half of the time the neuron does not fire (refer to Figure 3(d) and notice the large amount of even permutations with 00 activation). The model compensates for this by having multiple neurons dedicated to a single coset. Figure 5 shows how the unembedding layer aggregates four distinct $(H_{1},\;H_{1})$ circuits. For each of the neurons the effects on the final logits is simple: the product $\sigma_{l}\sigma_{r}$ is not in $H_{1}$, so the the unembedding uses the signal from these neurons to decrease the logits of all of the permutations in $H_{1}$ and increase the logits of all of the permutations that are in $H_{1}$. For a single neuron, however, there is non-trivial overlap between $H_{1}$ and not-$H_{1}$ at 00. The mistakes of each individual neuron cancel out when they are all added together, a single neuron may fail to activate for a permutation, but it is very unlikely that all four $(H_{1},\;H_{1})$ neurons will fail to fire for that neuron. As we see in Figure 5, with all four neurons there is minimal overlap in the logits.
@@ 6 The Process of Reverse Engineering
In this section, we recite the process of reverse engineering the coset circuit. First, we observe the activations and build up a theory for their mechanism. Second, we confirm the sufficiency and necessity of the circuits we have identified via ablations. Finally, we rigorously validate our hypothesis by testing the properties of the circuit through causal interchange interventions.
@@ 6.1 Identifying Coset Circuits
To do mechanistic interpretability you must spend some time staring at the weights. The group Fourier transform did, however, give us a place to start looking. Individual neurons would concentrate entirely on irreducible representations of $S_{n}$, the “frequencies” of the group Fourier transform. It was as if the activations of each neuron were a pure sine wave, as in [40]. Harmonic analysis over $S_{n}$ is subtler than in the classical case, however, and this was not enough for us to propose a concrete mechanism. After extensive manual effort and many dead-ends, we realized that a single neuron had its activations concentrated on the cosets of $F_{20}$, the Frobenius group of order 20 [14]. All of the activations for permutations in the same coset had the same value, with very little variance. This was so strange and unexpected that it warranted further investigation, which ultimately led us to discover the coset circuit.
For a function $f:G\rightarrow\mathbb{R}$, let $C_{H}(f)$ denote the the degree to which $f$ concentrates on the cosets $H\leq G$ where:
@@
Explicitly $C_{H}(f)$ calculates the degree to which restricting to the cosets of $H$ reduces the variance of $f$, $C_{H}(f)<1$ implies that $f$ can meaningfully be understood better by looking at the values that it takes on cosets. Since $S_{5}$ only has 156 subgroups  [^1] Verified with Sage [50] and GAP [20]., it is tractable to do an exhaustive search and calculate
@@
the subgroup that minimizes the variance of $f$ for every neuron in the model.
Figure 6: Ablation:Two cases of ablation. The full lines indicate ablating the double coset circuits in S5 and S6, the dashed lines indicate ablating the other neurons that don’t form the double coset circuits.
@@ 6.2 Ablations
To validate that the coset circuit is not only sufficient but also necessary to implement multiplication in $S_{n}$, we conduct ablations to observe the change in accuracy at inference time in two different ways: when the identified circuit is ablated and when the rest of the network is ablated.
If the identified circuit is necessary to solve the task, then we expect to see a drop in the accuracy when the identified circuit is removed from the network. If it is additionally the primary circuit used by the model, we expect the accuracy to remain high when we ablate the rest of the neurons and only keep the ones that form the coset circuits.
In figure 6, we observe a similar trend in both cases of ablations in both $S_{5}$ and $S_{6}$. We selected neurons that concentrate on one irrep above the threshold of x percent. When we only keep the neurons that have a concentration higher than 90%, the accuracy remains perfect in $S_{6}$ over 24 runs with various model sizes. The accuracy is near perfect (0.982) in $S_{5}$ over 107 runs. However, when we ablate these neurons, the accuracy drop to only 0.04 in S6 and 0.08 in $S_{5}$. This aligns with our initial hypothesis suggesting that the coset circuits are necessary and sufficient.
We also observe that the performance of $S_{5}$ (blue line) has a higher variance than $S_{6}$ (red line) We hypothesize that this is due to the smaller model sizes used in $S_{5}$.
@@ 6.3 Causal Interventions
Base Model 0.9999 1.97e-6
Embedding Swap 0.01 4.76
Switch Left and Right Sign 1 1.97e-6
Switch Left Permutation Sign 0 22.39
Switch Right Permutation Sign 0 22.36
Perturb $\mathcal{N}(0,0.1)$ 0.9999 2.96e-6
Perturb $\mathcal{N}(0,1)$ 0.978 0.0017
Absolute Value Non-Linearity 1 3.69e-13
Perturb $\mathcal{N}(1,1)$ 0.88 0.029
Perturb $\mathcal{N}(-1,1)$ 0.98 0.0021
To rigorously test the properties of the coset circuit, we carefully designed causal interchange interventions [21] targeting the properties of different components in the circuits. We need to confirm that our model of the circuit serves as a faithful causal abstraction for the behavior of the actual neural circuit. We aggregated runs over 128 $S_{5}$ models of different and recorded their average loss and accuracy. Initially, over the initial models without intervention, we have accuracy extremely close to 1.
@@
ReLU Clip
If the pre-activations are greater than zero and aren’t filtered by the ReLU, then the magnitude of the activations does not matter, there is no particular information in or distinction between a small or large pre-activation value. To test this, we patch the activations of neurons only when they are greater than some small constant and replace them with a normally distributed value centered around the mean activation of that neuron.
@@
7 Discussion
7.1 Communication and Replication in Mechanistic Interpretability
The project of mechanistic interpretability strives to understand and explain inner workings of a machine learned model. As the large models that are popular today are immensely complicated and difficult to deconstruct, a substantial amount of mechanistic interpretability research is dedicated to either small models or work that allows insights gleaned from small models to be extended to larger models.
Unfortunately, our analysis reveals that even allegedly simple cases can be very challenging to get right. By revising the problem considered by Chughtai et al. [6] and finding drastically different mechanisms after extensive and earnest effort, our work reveals how challenging it can be to get the full picture correct. We refer the reader to Appendix F for a full side-by-side comparison of our work with Chughtai et al. [6].
Based on our analysis of twelve mechanistic interpretability papers [6, 9, 11, 22, 24, 25, 37, 40, 41, 55, 58, 59] we have identified five core steps in mechanistic interpretability work:
  1. 1.
    Pick a reference data set or distribution that is correlated to the task.
  2. 2.
    Identify a subset of nodes in the model’s computation graph that acts as a neural circuit such that the activations of the circuit are causally connected to the reference distribution.
  3. 3.
    Establish that this circuit accounts for some substantial portion of the performance on the task.
  4. 4.
    Come up with a model of the circuit that faithfully abstracts the its functionality.
  5. 5.
    Causally test that the behavior of your model is faithful across data distributions.
Many new techniques for accomplishing step 2 such as activation and path patching [22, 53, 55] have been developed, though in our case we relied on the spectral properties of permutation groups to tell us where to stark looking. As [55] discuss, step 3 can also be difficult in larger models.
Steps 4 and 5 are a difficult, manual, and largely ad-hoc process. Once we have pinpointed the subset of neurons that is doing something important we still need to interpret what it is doing: what information those neurons are encoding, the precise mechanism with which those neurons act and are acted upon by the rest of the network, etc…
We believe that the root issue with [6] is that they could not find a precise mechanism for matrix-matrix multiplications in a 1-layer network. They got to step 4 and came up with a compelling theory that fit the evidence and tied in nicely with prior work such as [3, 40]. Without a concrete and detailed theory for how the model implements their proposed algorithm their ability to move on to step 5 and confirm that the network behaves as they predict is limited. Our work and that of [6] together are a case study in why performing detailed and circuit-specific tests is necessary. The spectral evidence in the weights and logits that both of our teams found did not uniquely determine a single circuit, but was consistent with both the GCR algorithm and the coset circuit. It was only after performing the causal interchange interventions detailed in Section 6.3 that we were confident that we understood what the model was doing.
We believe that one of the most promising paths forward for making step 5 as easy and rigorous as possible is extending the work of Geiger et al. [21]. There is a long history of interpretability techniques giving inconsistent and misleading evidence [1, 4, 13, 27], and recent results show that the most modern techniques can are vulnerable to similar issues [5, 18, 25, 34, 37]. There are ways to mitigate those weaknesses and in our review of the literature we were encouraged to see that most projects people went to great lengths to causally test specific features of the circuits that they propose under a variety of data distributions. Until we have developed better tools to make performing these tests easier, however, their application will remain ad hoc.
There is another subtler issue caused by the difficulty of communicating the precise mechanisms of neural circuits once they have been identified. Lieberum et al. [32] talk about the dilemma they faced when describing the circuits they found with pseudocode: they could remain faithful to the matrix algebra the circuit uses or provide an understandable, high-level explanation, but struggled to do both. In our work we have opted to keep our explanations as concrete and close to the actual neurons as possible, but doing so in a way that people will actually understand has been difficult. Effective communication of every minor detail and tweak to our architecture, such as concatenating the embeddings of the left and right permutations, is crucial for our audience to properly understand the circuit we are trying to explain. We need to develop better and more specific jargon, a grammar for communicating the circuits we find in a way that other researchers can understand. We know that the researchers who are doing the research can be fooled by explanations, but the situation for people reading and reviewing interpretability research is even worse. Our experience doing this project and writing this paper has led us to believe that developing a concise language for communicating our theories and the circuits we find is essential for facilitating progress in interpretability.
Finally, even simple examples can give rise to immense complexity. In Zhang et al. [58] the authors show that which algorithm a network learns can be sensitive to details of the experimental set-up such as the learning rate and initialization. While we were able to exactly replicate the results of Chughtai et al. [6] due to precise reporting, the release of their code, and personal communication, it is unfortunately common for machine learning researchers to not fully report their experimental set-up or release all of their evaluation artifacts. In such settings, it may be impossible to distinguish between different results due to genuinely different model behaviors versus ones due to erroneous analysis.
7.2 Spectral Information and Grokking
There is now a growing body of work making links between models in highly constrained domains that exhibit perfect generalization and the domain-appropriate Fourier transform. Nanda et al. [40] find that models trained on modular addition encode the dual representation of the cyclic group $C_{p}$ directly and Barak et al. [3] hypothesize that the model is using the spectral gap calculated via the Fourier transform on the Boolean cube inherent to the noisy sparse parities problem. We now find that a 1-hidden-layer network uses Fourier-sparse functions to solve multiplication on permutation groups. There are known results from complexity and computational learning theory that functions with low-degree Fourier concentration are, in a precise sense, “learnable” [42]. Though this still does not provide direct evidence of a link, we find it intriguing and an exciting direction for future research.
7.3 Duality and Interpretable Bases
Though the precise connections between the learning dynamics, the spectral characteristics of the weights, and the generalization properties of the model remain opaque, it is absolutely the case that this project would have been much more difficult (and perhaps impossible) without an interpretable basis from which to view the weights and activations.
The Fast Fourier Transform (FFT) [10] is widely regarded as one of the most important algorithms of all time. The generalizations of harmonic analysis to more complex geometries has also been integral to physics in the last century [2]. It is not surprising that similar algorithms would prove invaluable in machine learning and mechanistic interpretability.
There is also a rich representation theory of semigroups and monoids [51] and existing work describing FFTs for inverse semigroups [35, 36]. It appears promising that the surprising results of Liu et al. [33] might become concretely interpretable via the representation theory of the monoids that they study.
8 Conclusion
In this paper we performed an a circuit level analysis to discover the concrete mechanism a one layer fully connected network uses to solve group multiplication in $S_{5}$ and $S_{6}$. We showed that the model decomposes $S_{5}$ and $S_{6}$ into its cosets and uses that structural information to perfectly implement the task. We related how we found the circuit using the Group Fourier Transform and our methodology for confirming that our model of the circuit was faithful to its true behavior. Finally, we surveyed some modern research in mechanistic interpretability, their methodologies and the types of evidence they use to confirm their work. The causal interchange analysis that was crucial for us to confirm our hypotheses is difficult to perform and, to our knowledge, has not been applied to a real-world problem. This is an important obstacle for future research to overcome.

### v2

@@ 5.2 Conjugate Subgroup Circuit
All four ways to multiply two cosets of $A_{n}$ are well-defined. For each of the four options (even-even, odd-even, etc…) we know which coset of $A_{n}$ the product will be in, but no other subgroup of $S_{n}$ has this property. The model instead learns to use sets of conjugate subgroups. Recall that $H_{i}<S_{n}$ is the subgroup isomorphic to $S_{n-1}$ that fixes the element $i\in[n]$ in the $i^{\text{th}}$ place and $\tau_{ij}$ is the permutation that swaps $i$ and $j$. Any two $H_{i}$ and $H_{j}$ are conjugate to each other, $\tau_{ij}H_{i}\tau_{ij}=H_{j}$ and $\tau_{ij}H_{j}\tau_{ij}=H_{i}$. This means that there are two shared cosets between $H_{i}$ and $H_{j}$, because $H_{i}\tau_{ij}=\tau_{ij}H_{j}$ and $H_{j}\tau_{ij}=\tau_{ij}H_{i}$. The model implements the full group multiplication by picking out the shared cosets of conjugate subgroups.
As an example, consider a neuron that corresponds to $H_{1}$ for the left permutation and $H_{5}$ for the right permutation. The shared coset is $H_{1}\tau_{15}=\tau_{15}H_{5}$, the set of all $\sigma\in S_{5}$ with $\sigma(1)=5$. The pre-activations for the left and right permutations will be:
$L(\sigma)=\begin{cases}4\quad\sigma\in H_{1}\\ 2\quad\sigma\in H_{1}\tau_{12}\\ 0\quad\sigma\in H_{1}\tau_{13}\\ -2\quad\sigma\in H_{1}\tau_{14}\\ -4\quad\sigma\in H_{1}\tau_{15}\\ \end{cases}$ $R(\sigma)=\begin{cases}-4\quad\sigma\in\tau_{15}H_{5}\\ -2\quad\sigma\in\tau_{25}H_{5}\\ 0\quad\sigma\in\tau_{35}H_{5}\\ 2\quad\sigma\in\tau_{45}H_{5}\\ 4\quad\sigma\in H_{5}\\ \end{cases}$ (1)
The final activation is still $\texttt{ReLU}(L(\sigma_{l})+R(\sigma_{r}))$, but now there are twenty-five possible pairs of cosets. All twenty-five combinations can be boiled down to two meaningful cases:
  1. 1.
    If $L(\sigma_{l})+R(\sigma_{r})=0$, then $\sigma_{l}\sigma_{r}$ is in the shared coset $H_{1}\tau_{15}$.
  2. 2.
    If $L(\sigma_{l})+R(\sigma_{r})\neq 0$, then $\sigma_{l}\sigma_{r}$ is not in the shared coset $H_{1}\tau_{15}$.
Each left coset $yH_{5}$ has a paired right coset $H_{1}x$ such that $H_{1}xyH_{5}=H_{1}\tau_{15}=\tau_{15}H_{5}$. The discrete values that $L$ and $R$ can take are precisely tuned so that those pairs of left and right cosets cancel out. Just like with the sign neuron, information about the pre-activation being negative is lost with the ReLU. This lost information has to be made up with extra neurons that correspond to $(H_{1},H_{5})$ and assign different values to the cosets. For example, a different neuron that uses $-L(\sigma_{l})-R(\sigma_{r})$ will fail to fire for a different set of permutations. The combination
$\texttt{ReLU}(L(\sigma_{l})+R(\sigma_{r}))+\texttt{ReLU}(-L(\sigma_{l})-R(\sigma_{r}))$
will be much closer to a perfect on/off switch for coset membership.
@@ 5.3 Decoding Permutations with Coset Membership
[…]
@@ 6.1 Identifying Coset Circuits
[…]
@@
Where $\operatorname{Var}[f|_{gH}]$ is the variance of $f$ when the domain is restricted to the coset $gH$. Intuitively $C_{H}(f)$ calculates the degree to which restricting to the cosets of $H$ reduces the variance of $f$. If $C_{H}(f)<1$ it implies that the activations $f$ can meaningfully be understood better by looking at the values that it takes on the cosets of some subgroup. Recall that a single neuron is a function $N_{i}:S_{n}\times S_{n}\rightarrow\mathbb{R}$ is the sum of two functions $G\rightarrow\mathbb{R}$, one for the left and right permutations, respectively. We can calculate $\min C_{H}$ for each. Take as an example $N^{l}_{115}$, the neuron shown in Figure 3. At 100,000 steps (on the far left) $\operatorname{Var}[N^{l}_{115}]=5.23$. Its activations are not concentrated on the specific cosets of $F_{20}$, however, and $C_{F_{20}}(N^{l}_{115})=2.96$. At 130,000 steps (on the far right) $\operatorname{Var}[N^{l}_{115}]$ has increased to $9.06$, but $C_{F_{20}}(N^{l}_{115})<10^{-5}$. The distribution within each coset of $F_{20}$ has close to zero variance.
We see a typical example of what this looks like for the entire model in Fig. 4. As the validation loss approaches a small value, there is a rapid transition from the median coset concentration being approximately $1$, to a minuscule value.
Even if it is apparent that a neuron is taking on discrete values and is a good candidate for being a coset neuron, it is difficult to tell by sight which subgroup the neuron is activating for. $S_{5}$ and $S_{6}$ only have 156 and 1,455 subgroups, respectively,  [^3] Sequence A005432 OEIS [42] so it is tractable to do an exhaustive search and calculate
@@
the subgroup that minimizes the variance of $f$ for every neuron in the model. Running these calculations shows that for the 128 $S_{5}$ models and 100 $S_{6}$ models we trained over 99.2% of the neurons in the linear layer had $\min_{H\in\operatorname{Sub}(G)}C_{H}(f)<1.0$, and the vast majority of those were less than $10^{-6}$.
With the ability to calculate directly which neurons corresponded to which subgroup, our theories for exactly what the neurons were representing fell into place. The next step was to confirm that these neurons were actually responsible for the models’ performance.
Figure 4: The paired evolution of the the validation loss and $\min_{H\in Sub(H)}C_{H}$, which encodes the formation of coset circuits. Displayed is the $S_{5}$ model with random seed 1. Different runs will form coset circuits at different times in training, but the effect is representative.
Figure 5: We perform ablations by re-calculating the accuracy after removing any neurons $N_{i}$ that have $\min_{H\in\operatorname{Sub}(G)}C_{H}(N_{i})$ greater than (top figure) or less than (bottom figure) the thresholds on the x-axis.
@@ 6.2 Ablations
We have described how coset neurons function and how they can be identified. We will now show via ablations that coset neurons are not solely sufficient but also necessary to implement multiplication in $S_{n}$. We conduct ablations by removing neurons which have a coset concentration $\min_{H\in\operatorname{Sub}(G)}C_{H}(N_{i})$ above a threshold.
If coset circuits are in fact responsible for the performance of our models, then we expect to see no change in the accuracy when the neurons that have not converged onto the cosets of a subgroup are removed from the model. This is precisely what we see on the far right of Figure 5. Of the 128 $S_{5}$ models that we trained, 126 models saw no change in the accuracy when we removed the neurons with $\min_{H\in\operatorname{Sub}(G)}C_{H}(N_{i})\geq 1$ (the far right of Figure 5). Recall that if $C_{H}(N_{i})\geq 1$, restricting to the cosets of $H$ at best does not change the variance of $f$. Of the two models that did show a decrease in accuracy, they decreased to 99% and 98%.
We see more between-run variation when we remove more neurons. The median model has 24 out of 128 neurons with $\min C_{H}(N_{i})\geq 10^{-5}$, but the $50^{\text{th}}$ and $25^{\text{th}}$ percentile accuracy is still 100%. It is not until we set the threshold to $10^{-6}$ that the $25^{\text{th}}$ percentile moves at all. When we set the threshold at $10^{-7}$ the performance for many models has collapsed, but the median model has had 42 neurons removed and the median accuracy is still 100%. Recall also that the neuron shown in the far right of Figure 3 has a coset concentration of $10^{-5}$.
The overwhelming majority of neurons are identifiable as coset neurons. Of those neurons, those with the very highest concentration on cosets account for the largest portion of each model’s performance.
@@ 6.3 Causal Interventions
Base Model 99.99% 1.97e-6
Embedding Swap 1% 4.76
Switch Left and Right Sign 100% 1.97e-6
Switch Left Permutation Sign 0% 22.39
Switch Right Permutation Sign 0% 22.36
Perturb $\mathcal{N}(0,0.1)$ 99.99% 2.96e-6
Perturb $\mathcal{N}(0,1)$ 97.8% 0.0017
Absolute Value Non-Linearity 100% 3.69e-13
Perturb $\mathcal{N}(1,1)$ 88% 0.029
Perturb $\mathcal{N}(-1,1)$ 98% 0.0021
To rigorously test the properties of the coset circuit, we carefully designed causal experiments to test specific properties of in the circuits. We observe a circuit’s behavior over the entire data distribution (the full group $S_{n}$) and we see that our model of the circuit is consistent with the behavior of the true circuit. To confirm that our model of the circuit is correct, however, we need to “break” the circuit in targeted ways and test that it behaves in the way we predict. Neural circuits are complex enough that observational evidence is not enough. We aggregated runs over 128 $S_{5}$ models of different and recorded their average loss and accuracy. Initially, over the initial models without intervention, we have accuracy extremely close to 1.
@@
7 The Group Composition via Representations Algorithm
Our experimental setup is identical to that of Chughtai et al. [4], but our analysis led us to a different conclusion.Chughtai et al. [4] proposed the “Group Composition via Representations” (GCR) algorithm. They show that, given an irrep $\rho$ of $S_{n}$, $\operatorname{argmax}_{c\in S_{n}}\operatorname{tr}[\rho(a)\rho(b)\rho(^{-1}c)]=ab$ and propose that this is the algorithm the model is implementing. This requires that not only store the matrix irreps, but that the model perform the matrix multiplication within its mechanism. We find that most of the evidence [4] put forward is also consistent with coset circuits. The other evidence we were not able to independently replicate. We also find evidence that, to our understanding, is not consistent with the GCR algorithm but is explained by coset circuits.
7.1 Our Interpretation of the Evidence for GCR
Chughtai et al. [4] put forward four main pieces of evidence, which we restate here for clarity: (1) Correlation between the model’s logits and characters of a learned representation $\rho$. (2) The embedding and unembedding layers function as a “lookup table” for the representations of the input elements $\rho(a),\;\rho(b)$ and the inverse of the target $\rho(c^{-1})$. (3) The neurons in the linear layer calculate the matrix product $\rho(a)\rho(b)=\rho(ab)$. (4) Ablations showing that the circuit they identify is responsible for the majority of the model’s performance. Many of these points are equally consistent with the coset circuit and the other we could not find evidence for.
Irrep Look Up Table We were not able to find any evidence that the embedding or unembedding layers function as a look-up table for any representation except for the one-dimensional sign representation. We did find that the model’s weights and activations concentrate on specific irreps in the group Fourier basis. This is due, however, to concentration on cosets of specific subgroups, not because the matrix representations are realized anywhere in the weights. The relationship between functions that are constant on cosets and specific irreps is shown in Appendix G.2.
Logit Attribution The trace of a group representation is referred to as the “character” and often denoted $\chi$. We find that the model’s logits correlate with the character $\chi_{\rho}(abc^{-1})$ when the irrep $\rho$ appears in the Fourier transform of the model’s weights. This is not, however, because the model has implemented the matrix product $\rho(ab)\rho(c^{-1})$, but because the model is “counting” the number of cosets that $ab$ and $c$ are both in. We prove in G.2, if the cosets are of conjugate subgroups that have their Fourier transform concentrated on the irrep $\rho$ (as we observe for the models in question), then the number of shared cosets will also correlate with the characters of $\rho$.
Matrix Multiplication of Irreps We were not able to find any evidence that the linear layer implements matrix multiplication, again excluding scalar multiplication of the sign irrep.
7.2 Evidence GCR Does Not Explain
The Difference Between Subgroups and Irreps The GCR algorithm and coset circuit cannot be equivalent because there is not, in fact, a one-to-one relationship between cosets and irreps. Most subgroups of $S_{n}$ have their Fourier transforms concentrate on more than a single group (see Table 5 for the spectral properties of all of the subgroups of $S_{5}$), indeed this needs to be the case as there are many more subgroups than irreps. Please refer to Table 4 for a concrete comparison and Appendix 7.2 for an asymptotic analysis. We also observe coset circuits for some subgroups such as $D_{10}$ [^4] The dihedral group of order 10, the symmetry group of a pentagon. will have coset circuits concentrated on both (3, 2)32(3,\;2)( 3 , 2 ) or (2, 2, 1)221(2,\;2,\;1)( 2 , 2 , 1 ), depending on the run. The GCR algorithm would treat these as different circuits, though their behavior is in fact identical.
Unembedding Correlations of Neurons We observe that the correlation between in the unembedding of neurons that concentrate on the same coset is on average $81.4\%$ (see Table 3). The correlation between neurons concentrated only on the same conjugacy class of subgroup (e.g. $H_{1}$ and $H_{2}$) is on average $-0.2\%$. The neurons that represent subgroups in the same conjugacy class will oftentimes, though not always, be concentrated on the same irrep. The model is treating cosets together but the irreps and conjugacy classes separately.
The concentration of the model’s activations on irreps of $S_{n}$ is striking evidence and the GCR algorithm that [4] detail could indeed solve the problem of group multiplication. The coset circuit is also consistent with all of the evidence that [4] provide and is additionally consistent with evidence that the GCR algorithm does not explain.
8 Discussion and Conclusion
We performed a circuit level analysis to discover the concrete mechanism a one layer fully connected network uses to solve group multiplication in $S_{5}$ and $S_{6}$. We showed that the model decomposes $S_{5}$ and $S_{6}$ into its cosets and uses this structural information to perfectly implement the task.
Though our work concerns a toy problem, we highlight a core takeaway that applies broadly to the field of interpretability: we must treat proposed neural mechanisms as theories until they have been thoroughly tested.
When we identify what we believe to be a circuit within a larger network found via techniques such as [7, 19], we have taken the first step towards mechanistically understanding how a model performs a task. The evidence we have for the circuit’s role in that task is, however, fundamentally observational and correlational. The nodes in the circuit’s computation graph are causally connected, but the relationship between the action of those nodes is only observed to be correlated to a certain task with respect to a distribution. This is valuable information to have, but the understanding that it imparts is limited and must be recognized as such.
When beginning this project we quickly noticed that the activations of sub-circuits of our model were concentrated on specific irreps of $S_{n}$. It was only with additional investigation that we were able to attach semantic meaning to this phenomenon. We observed that the neurons concentrated on a single irrep were activating for specific subgroups. The hypothesis of the coset circuits had formed, but it was still only a theory. The facts we had observed were incontrovertible, but their reason was unclear. It was only after performing the causal experiments detailed in Section 6.3 that we became confident we understood the mechanism. The simple reality is that more than one theory can be consistent with observational data, especially when that data only comes from a small subset of the full distribution. There is a long history of scholarship showing that interpretability techniques, including state-of-the-art, can give be misleading and contradictory results [1, 2, 3, 9, 14, 23, 25, 34, 35].
In doing this work we had many advantages not available when interpreting real-world models: access to the entire distribution, an orthonormal basis for the function space of the network, and a relatively small model. The task of multiplication in $S_{n}$ is deterministic and very well studied, we had many mathematical tools to bring to bear in analyzing the model. Even still, this project was quite challenging and the circuits we found surprised us. Interpreting real models will be even difficult. We encourage future work to apply interpretability tools cautiously and validate observational results with rigorous experimental tests.
Impact Statement
This paper presents work whose goal is to make the function and mechanisms of deep neural networks interpretable to humans. We present methods for reasoning about counterfactual and out of distribution behavior in the models that we train. Though our setting is too small to be directly relevant to real-world use cases, we hope that similar techniques will be able to test, audit, and monitor deep neural networks that have been deployed in the real world. We also present results that urge caution and humility when attempting to interpret neural networks. We believe that robust and effective interpretability techniques may mitigate some societal harms that could arise from the use of deep neural networks, but that mistakenly trusting illusory interpretability techniques could be disastrous.